if 15 bales of hay feeds 10 horeses, how many will 18 bales of hay feed?
Found 2 solutions by Maths68, stanbon:
Answer by Maths68(1474)

15 bales of hay feeds................. 10 horses
15/15 bales of hay feeds................. 10/15 horses
1 bale of hay feeds................. 2/3 horses
1*18 bales of hay feeds.................18*2/3 horses
18 bales of hay feeds.................6*2 horses
18 bales of hay feeds.................12 horses

Answer by stanbon(75887)

if 15 bales of hay feeds 10 horeses, how many will 18 bales of hay feed?
----
# of bales and # of horses are directly related:
h = k*b
Solve for "k" usiing "15 bales feeds 10 horses"
10 = k*15
k = 2/3
---
Equation:
h = (2/3)b
----
how many will 18 bales of hay feed?
h(18) = (2/3)18
h(18) = 12 (# of horses fed by 18 bales)
================
Cheers,
Stan H.
====================